OpenStudy (awesomedude2078):

Calculate the paychecks of the following people. Select the total hours and paycheck. Total Hours Rate Paycheck M T W T F Sue Black 8 8 8 8 8 $6.75 Jerry Atrix 7 ¼ 3 6 ¾ 8 5 $5.15 Sam Ting 7 ¾ 7 ¾ 7 ¾ 7 ¾ 7 ¾ $8.75

563blackghost (563blackghost):

It shows on the table that they have a `rate` of (blank) dollars that would mean that they work that many dollars an hour.... It is also showing us how many hours they worked each day....You would simply add up all those hours in which it would give you the total hours for the week.... To find the paycheck of the week you would multiply the rate by the total hours of the week...

563blackghost (563blackghost):

So lets do Sue Black we see that she worked 8 hours for Monday-Friday...to find how many hours she worked we would add up all the hours.... \(\huge{8+8+8+8+8 = 8 \times 5 = 40 ~hours}\) So we know she worked `40 hours`.....Now we see that she worked at a rate of `$6.75` (if im reading this correctly) that would mean that we would multiply the rate by the total hours so.... \(\huge\color{green}{40 \times 6.75=?}\)

OpenStudy (awesomedude2078):

270

563blackghost (563blackghost):

Correct :)

563blackghost (563blackghost):

Now you would follow the same process by adding up the total hours and multiplying it by their individual rates....This will give you the paycheck for each person as well as total hours...

563blackghost (563blackghost):

(by the way i love your prof pic :D)

OpenStudy (awesomedude2078):

thanks

OpenStudy (awesomedude2078):

yours is good too

563blackghost (563blackghost):

Thank you :D

OpenStudy (awesomedude2078):

you're welcome

563blackghost (563blackghost):

Do you understand the process I typed out?